Optical Channel Performance





September 2000

The optical chip is a Loral 2048 x 2048 CCD with a scale of 0.3 arcsec/pixel and about a 10' x 10' field of view. Due to amplifier problems, only half of the chip has been working in the past. During the last engineering run (September 6-9 2000) this problem has been fixed. Flat-fielding and photometric tests suggest that both halves of the CCD now work about equally well. Note that there are a few differences between the two halves: the left (new) half has a different bias level than the right (old) half (150 vs. 100 DN); the left half has a small nonfixed pattern present in the bias (amplitude ~4 DN) that is not seen in the right half; the first ~10 data columns of the left half have a reset-anomaly that causes very high counts in those columns (so high that no useful information is present). Otherwise the entire CCD seems to be working.

The gain was also reset on the CCD. The gain is now roughly 3.5 electrons/DN. Because the gain is set by the dwell time spent measuring the signal in each pixel, this also lowered the read noise to roughly 11 electrons. A consequence is that the CCD now takes about 20 seconds longer to read out than before.

The number of overscan columns were also reset. There are now 32 columns of overscan on each side of the CCD. There are, in addition, 16 columns of what we call "underscan". These are functionally the equivalent of overscan, although we suggest not using them to determine the bias level in the CCD.

Please consult the Ohio State YALO Telescope & Instrument News wepage for more details on YALO data reduction.

January 2000

During the latest engineering time in early January 2000, the following Andicam servicing and testing was performed:
The "wide R" filter was changed to a more standard "R" filter supplied by Nick Suntzeff.

While the instrument was off the telescope, Max Boccas, Eric Rubenstein, and Rene Mendez performed some tests of the telescope optical system with a direct CCD. Max adjusted the primary supports and reported 1" unguided images in a 10 minute exposure.

Darren Depoy re-aligned the IR and CCD channel and re-focussed the IR channel. The CCD images were typically 1" and the IR images slightly better during the focussing. As predicted, the range of positions for the IR focus mirror is very large (i.e. there are very good IR images over a large range of mirror positions). There is no need to re-focus the IR channel relative to the CCD channel with any filter combinations.